# DMRetriever: A Family of Models for Improved Text Retrieval in Disaster Management

### v1

@@ 4.1 DisastIR-DevLite and DisastIR-Test Construction
Prior works (Boteva et al., 2016; Yang et al., 2018; Thakur et al., 2021) typically construct validation sets by randomly splitting queries while keeping the passage corpus identical across validation and test sets, which makes repeatedly encoding the full passage corpus computationally expensive.
@@ 4.2 Baselines and Metrics
Baselines are drawn from two sources: (1) models evaluated in DisastIR, and (2) the latest high-performing open-source models on the MTEB retrieval benchmark  [^5] As of August 15, 2025. (Muennighoff et al., 2022) that have not been evaluated on DisastIR.
@@ 5.3 Ablations on Data Refinement Pipeline (RQ2)
Figures 3 and 7 highlight the importance of disaster management-specific and general-domain corpora in shaping model performance. General-domain data introduces diverse language patterns and semantic relevance, enhancing model robustness and adaptability (Xu et al., 2024a).
Disaster management corpus grounds model in the context of this domain, enhancing its ability to adapt to domain-specific scenarios. This yields substantial improvements, particularly on search intents that demand intensive domain knowledge (e.g., NLI and Twitter) (Yin et al., 2025), with gains of up to 2.9 points on Twitter in 596M variant. These results underscore the importance of disaster management-specific data for effective domain adaptation and model performance.
@@ 5.4 Ablations on Three-stage Training Frameworks (RQ3)
We also find that adding intent-specific instructions to query significantly improves performance (Figure 5c), with gains of up to 2.3 points for 109M variant. This suggests that instruction tuning helps model better align its embeddings with diverse search intents (Asai et al., 2022; Wang et al., 2023).
@@
Adaptation to bidirectional attention is vital for decoder-only backbones, yielding a significant performance improvement (Figure 5b) with up to 2.7 points improvement (596M variant). This indicates that bidirectional attention could yield more effective embeddings than causal attention, as it allows model to capture knowledge from future tokens (Lee et al., 2024a).
Contrastive pre-training is crucial for effective domain knowledge adaptation, especially for smaller models, yielding robust initialization for retrieval tasks (Figure 5c). The performance gains, however, decrease with increasing model size, as larger models may already possess the capacity to capture domain knowledge (Xu et al., 2024b).
@@ 6 Conclusion
A potential issue is test set contamination, where some test samples may overlap with training data (Sainz et al., 2023). We conduct detailed data contamination analysis in Appendix C, where we do not observe any overlap between train and test sets.
@@ Appendix A Details of DM-MTP Construction
For web crawling, we construct queries in the format “event type name + location name + .pdf”. We specifically target PDFs because they are more likely to contain structured, information-rich, and credible content, often originating from peer-reviewed publications or official institutions. To ensure coverage, we consider 301 disaster event types as defined in UNDRR (2020) and combine them with the names of the 50 U.S. states, given both the availability of high-quality resources in the U.S. context and the fact that our model is designed for English-language data.
All collected PDF files are then processed into semantically coherent text chunks through a multi-step pipeline proposed by Yin et al. (2025): (1) exact-URL deduplication, where duplicate documents are removed by identifying identical download links; (2) text extraction and preprocessing, where each PDF is converted into plain text and non-textual elements such as tables and figures are removed; (3) locality-sensitive hashing (LSH) deduplication, which eliminates near-duplicate documents with overlapping content; (4) semantic chunking, where cleaned documents are segmented into passages of fewer than 256 tokens to balance retrievability and semantic integrity; and (5) embedding-based near deduplication, where dense embeddings are computed for all chunks, an ANN index is built, and pairs with cosine similarity above 0.9 are discarded. Chunks obtained are regarded as disaster management passages in DM-MTP.
In the second and third stages, given a passage and the definition of a search intent, an LLM generates information-need statement. The information statement together with the passage is used to construct a user query and its directly relevant positive passage. Prompts for query generation and relevant passage generation based on disaster management-related passages under different search intents are adapted from Yin et al. (2025).
@@ Appendix C Data Contamination Analysis
Following Wang et al. (2023), we conduct test set data contamination analysis from two aspects: query overlap and passage corpus overlap.
@@
Following Wang et al. (2023), we perform a string-matching analysis between queries in the DisastIR-Test and those in the training sets (MTP and MTT), and find no overlap between training and test queries.
@@
While sharing the same passage corpus (e.g., DBPedia, NQ, and TriviaQA all using Wikipedia) is a standard evaluation practice in information retrieval and is not regarded as contamination (Wang et al., 2023), we adopt a stricter criterion by performing a string-matching analysis between passages in DisastIR-Test and those in MTP and MTT, finding no overlap between the training and test corpora.
@@ Appendix D Determination of $\alpha$ Value during Progressive Fine-tuning
We follow the ablation study results of Moreira et al. (2024) and set $\alpha=0.95$ for DMRetriever-596M, 4B, and 7.6B.
@@ Appendix G Detailed Reasons for Excluding Knowledge Distilled (KD) Baselines
For baselines (Wang et al., 2022; Awasthy et al., 2025) that use knowledge distillation (KD) during fine-tuning, a cross-encoder is typically employed as the teacher model. This process introduces substantial external knowledge, as the teacher cross-encoder has a more powerful architecture. Unlike bi-encoders (our primary baselines and DMRetriever), cross-encoders process queries and documents jointly, allowing deep token-level interactions that produce highly accurate relevance scores. These fine-grained scores are then used as soft labels for the student model, offering a much richer and more informative training signal than the sparse binary hard labels available to other models.
Consequently, student model does not learn from ground-truth data alone, but is guided by teacher’s high-quality interpretation of that data. For this reason, we exclude all KD-based baselines, which include E5-small, base, large-v2 (Wang et al., 2022), and granite-embedding-125m-english (Awasthy et al., 2025), to ensure fair comparisons.
@@ Appendix H Information on Baseline Models and Their Implementation
We adopt baseline models evaluated in DisastIR and additionally include recent models from MTEB that were not evaluated in DisastIR. They are categorized as four scales based on model parameter sizes: small (<=109M), medium (137M - 335M), large (434M- 1.5B), and extra large (XL) (>= 4B). Detailed specifications of the selected models are provided in Table 11, while their HuggingFace links and licenses are listed in Table 12.
@@ Appendix I Implementation Details of DMRetriever Training
For unsupervised contrastive pre-training, each model is trained for 1 epoch. We use 3 H200 GPUs for encoder-only backbones and 4 H200 GPUs for decoder-only backbones. To optimize GPU memory consumption, we adopt DeepSpeed ZeRO Stage 1 (Rasley et al., 2020) and gradient checkpointing. Gradient accumulation is set to 1, and the per-GPU batch size is maximized to increase the diversity of in-batch negatives (Xiao et al., 2023), leading to the global batch sizes reported in Table 7.
@@
Across all experiments, we use brain floating point (bfloat16) quantization and truncate input sequences to 512 tokens. The temperature value is set as 0.01. For each query, we set one positive passage and nine passages during pre-training and fine-tuning. For decoder-only backbones, we apply LoRA with rank $r=16$ and $\alpha=32$, following prior work (Lee et al., 2024a; Wang et al., 2023; Xu et al., 2024b), leaving hyperparameter tuning for future exploration.
@@
To implement the CBF method, we follow the methodology of Wang et al. (2022). Specifically, for each search intent, we use its best-performing retrieval model identified from the DisastIR benchmark as a reference (Table 4 for best-performing retrieval models). A synthetic pair is then discarded if its LLM-generated passage is not ranked within the top-2 results retrieved by this reference model.
@@
Dataset Size Task Link
GD-MTP:
Squad 87.6K QA https://huggingface.co/datasets/sentence-transformers/squad
Paq 4M* QA https://huggingface.co/datasets/sentence-transformers/paq
MSMARCO 1M QA https://huggingface.co/datasets/microsoft/ms_marco
Gooaq 3M* QA https://huggingface.co/datasets/sentence-transformers/gooaq
Eli5 325k QA https://huggingface.co/datasets/sentence-transformers/eli5
Nfcorpus 134K QAdoc https://huggingface.co/datasets/mteb/nfcorpus
MSMARCO-Document-Ranking 250k* QAdoc https://microsoft.github.io/msmarco/Datasets
Hotpotqa 170k QAdoc https://huggingface.co/datasets/BeIR/hotpotqa
Signal1m-generated-queries 1M* Twitter https://huggingface.co/datasets/BeIR/signal1m-generated-queries
Fever 250k* FactCheck https://huggingface.co/datasets/mteb/fever
Quora 15k STS https://huggingface.co/datasets/mteb/quora
Quora-duplicates 950k* STS https://huggingface.co/datasets/sentence-transformers/quora-duplicates
Altlex 113k STS https://huggingface.co/datasets/sentence-transformers/altlex
XNLI 133k* NLI https://huggingface.co/datasets/mteb/xnli
All-NLI 321k* NLI https://huggingface.co/datasets/sentence-transformers/all-nli
DM-MTP 3.29M
Total 15M
Table 9: Composition of the MTP dataset used for unsupervised contrastive pre-training in Stage 2. * indicates that only a subset of the original dataset is used.

### v2

@@ 4.1 DisastIR-DevLite and DisastIR-Test Construction
Prior works (boteva2016full; yang2018hotpotqa; thakur2021beir) typically construct validation sets by randomly splitting queries while keeping the passage corpus identical across validation and test sets, which makes repeatedly encoding the full passage corpus computationally expensive.
@@ 4.2 Baselines and Metrics
Baselines are drawn from two sources: (1) models evaluated in DisastIR, and (2) the latest high-performing open-source models on the MTEB retrieval benchmark  [^5] As of August 15, 2025. (muennighoff2022mteb) that have not been evaluated on DisastIR.
@@ 5.3 Ablations on Data Refinement Pipeline (RQ2)
Figures 3 and 7 highlight the importance of disaster management-specific and general-domain corpora in shaping model performance. General-domain data introduces diverse language patterns and semantic relevance, enhancing model robustness and adaptability (xu2024simrag).
Disaster management corpus grounds model in the context of this domain, enhancing its ability to adapt to domain-specific scenarios.
This yields substantial improvements, particularly on search intents that demand intensive domain knowledge (e.g., NLI and Twitter) (yin2025disastir), with gains of up to 2.9 points on Twitter in 596M variant. These results underscore the importance of disaster management-specific data for effective domain adaptation and model performance.
@@ 5.4 Ablations on Three-stage Training Frameworks (RQ3)
We also find that adding intent-specific instructions to query significantly improves performance (Figure 5c), with gains of up to 2.3 points for 109M variant. This suggests that instruction tuning helps model better align its embeddings with diverse search intents (asai2022task; wang2023improving).
@@
Adaptation to bidirectional attention is vital for decoder-only backbones, yielding a significant performance improvement (Figure 5b) with up to 2.7 points improvement (596M variant). This indicates that bidirectional attention could yield more effective embeddings than causal attention, as it allows model to capture knowledge from future tokens (lee2024nv).
Contrastive pre-training is crucial for effective domain knowledge adaptation, especially for smaller models, yielding robust initialization for retrieval tasks (Figure 5c). The performance gains, however, decrease with increasing model size, as larger models may already possess the capacity to capture domain knowledge (xu2024bmretriever).
@@ 6 Conclusion
A potential issue is test set contamination, where some test samples may overlap with training data (sainz2023nlp). We conduct detailed data contamination analysis in Appendix C, where we do not observe any overlap between train and test sets.
@@ Acknowledgments
Input:
Backbone $\mathcal{M}$ (encoder- or decoder-only); MLM training data $\mathcal{D}_{MLM}$; Difficulty level schedule $\{\alpha_{t}\}$ for progressive fine-tuning
Output: Final DMRetriever model
1 Stage 1: Bidirectional attention adaptation (decoder-only backbone)
2 if $\mathcal{M}$ is decoder-only then
3   Replace causal mask with full attention mask;
4   Train $\mathcal{M}$ on $\mathcal{D}_{MLM}$ using the MLM objective
5Stage 2: Unsupervised contrastive pre-training
6 Construct $\mathrm{MTP}$ via Algorithm 2;
7 Pre-train $\mathcal{M}$ on $\mathrm{MTP}$ using InfoNCE loss (Eq. 2) with in-batch negatives
8 Stage 3: Difficulty-aware Progressive Supervised Instruction Fine-tuning
9 foreach iteration $t$ do
10   Construct $\mathrm{MTT}\text{-}\alpha_{t}$ via Algorithm 3 with difficulty level $\alpha_{t}$;
11   Fine-tune $\mathcal{M}$ on $\mathrm{MTT}\text{-}\alpha_{t}$ using supervised InfoNCE loss (Eq. 4);
12   Update $\mathcal{M}$ with the best validation checkpoint
return $\mathcal{M}$
Algorithm 1 Three-stage training pipeline of DMRetriever
Input:
$\mathcal{C}_{GD}$: general-domain corpora; $\mathcal{R}_{dm}$: raw disaster-management documents (e.g., PDFs); $\mathcal{I}$: search intents; $\Gamma_{i}$: heuristic rule for intent $i$; $\mathcal{E}^{need}_{i},\mathcal{E}^{query}_{i}$: in-context examples for information-need generation, and for query–positive passage generation;
$\mathrm{Prompt}^{need},\mathrm{Prompt}^{query}$: prompt templates for information-need generation, and for query–positive passage generation
Output:
$\mathrm{MTP}=\mathrm{GD\mbox{-}MTP}\cup\mathrm{DM\mbox{-}MTP}$
1
21. Build DM-MTP:
3 foreach $d\in\mathcal{R}_{dm}$ do
4 $\mathcal{C}_{dm}\leftarrow\mathrm{ProcessToPassages}(d)$;
5
6foreach $p\in\mathcal{C}_{dm}$ do
7 $s\leftarrow\mathrm{LLM}(p,\mathcal{I},\mathcal{E}^{need}_{i},\mathrm{Prompt}^{need})$;
8 $(q,p^{+})\leftarrow\mathrm{LLM}(s,p,\mathcal{E}^{query}_{i},\mathrm{Prompt}^{query})$;
9 $\mathrm{DM\mbox{-}MTP}\leftarrow\mathrm{DM\mbox{-}MTP}\cup\{(q,p^{+})\}$;
10
11
122. Build GD-MTP:
13 foreach $r\in\mathcal{C}_{GD}$ do
14   Identify intent $i$ (by dataset design);
15 $(q,p^{+})\leftarrow\mathrm{ApplyHeuristic}(r,\Gamma_{i})$;
16 $\mathrm{GD\mbox{-}MTP}\leftarrow\mathrm{GD\mbox{-}MTP}\cup\{(q,p^{+})\}$;
17
18
193. Combine and output:
$\mathrm{MTP}\leftarrow\mathrm{GD\mbox{-}MTP}\cup\mathrm{DM\mbox{-}MTP}$
Algorithm 2 Construction of MTP
QA QAdoc Twitter FactCheck NLI STS
Top 1 NV-Embed-v2 Linq-Emb-Mis SFR-Emb-Mis Linq-Emb-Mis NV-Embed-v2 NV-Embed-v2
Top 2 Linq-Emb-Mis NV-Embed-v2 inf-retriever-v1 SFR-Emb-Mis gte-Qwen2-1.5B-ins inf-retriever-v1
Top 3 inf-retriever-v1 SFR-Emb-Mis Linq-Emb-Mis NV-Embed-v2 inf-retriever-v1-1.5b gte-Qwen2-7B-ins
Table 4: Top-3 IR models across different search intents identified in DisastIR. Linq-Emb-Mis and SFR-Emb-Mis are shorten for Linq-Embed-Mistral and SFR-Embedding-Mistral
Task Instruction Format
QA Given the question, retrieve most relevant passage that best answers the question
QAdoc Given the question, retrieve most relevant document that answers the question
Twitter Given the user query, retrieve the most relevant Twitter text that meets the request
FactCheck Given the claim, retrieve most relevant document that supports or refutes the claim
NLI Given the premise, retrieve most relevant hypothesis that is entailed by the premise
STS Given the sentence, retrieve the sentence with the same meaning
Table 5: Instruction formats used for each search intent.
Input:
$\mathcal{Q}_{GD}$: queries from GD-MTP (with $(q,p^{+})$ pairs); $\mathcal{Q}_{DM}$: queries from DM-MTP; NN: top-NN passages for mutual-agreement; $\mathcal{M}_{top3}$, $\mathcal{M}_{best}$: top-3 and best IR models per intent (from DisastIR); $\alpha$: difficulty level for hard negative selection
Output: $\mathrm{MTT}\text{-}\alpha=\mathrm{GD\mbox{-}MTT}\text{-}\alpha\cup\mathrm{DM\mbox{-}MTT}\text{-}\alpha$
1
21. Mutual-agreement false positive filter
3 foreach $q\in\mathcal{Q}_{DM}$ do
4   Identify intent $i$;
5 foreach $m\in\mathcal{M}_{top3}(i)$ do
6      Retrieve top-NN passages $R_{m}(q)$;
7
8 $\mathcal{C}_{q}\leftarrow\bigcap_{m\in\mathcal{M}_{top3}(i)}R_{m}(q)$;
9 if $\mathcal{C}_{q}\neq\emptyset$ then
10      Add $(q,\text{top-1}(\mathcal{C}_{q}))$ to $\mathcal{D}^{+}_{DM}$;
11
12
[…]
14
152. Difficulty-aware hard negative mining
16 foreach $(q,p^{+})\in\mathcal{D}^{+}$ do
17   Identify intent $i$;
18   Retrieve top-200 $R_{best}(q)$ using $\mathcal{M}_{best}(i)$;
19 $\mathcal{H}_{q}\leftarrow\{p^{-}\in R_{best}(q)\setminus\{p^{+}\}\mid\mathrm{sim}(q,p^{-})<\mathrm{sim}(q,p^{+})\times\alpha\}$;
20   Add $(q,p^{+},\text{top-}k(\mathcal{H}_{q}))$ to $\mathrm{DM\mbox{-}MTT}\text{-}\alpha$ if $q\in\mathcal{Q}_{DM}$ else $\mathrm{GD\mbox{-}MTT}\text{-}\alpha$;
21
22
233. Combine and output: $\mathrm{MTT}\text{-}\alpha\leftarrow\mathrm{GD\mbox{-}MTT}\text{-}\alpha\cup\mathrm{DM\mbox{-}MTT}\text{-}\alpha$;
return $\mathrm{MTT}\text{-}\alpha$
Algorithm 3 Construction of MTT-$\alpha$
@@ Appendix A Details of DM-MTP Construction
For web crawling, we construct queries in the format “event type name + location name + .pdf”. We specifically target PDFs because they are more likely to contain structured, information-rich, and credible content, often originating from peer-reviewed publications or official institutions. To ensure coverage, we consider 301 disaster event types as defined in undrr2020hazard and combine them with the names of the 50 U.S. states, given both the availability of high-quality resources in the U.S. context and the fact that our model is designed for English-language data.
All collected PDF files are then processed into semantically coherent text chunks through a multi-step pipeline proposed by yin2025disastir: (1) exact-URL deduplication, where duplicate documents are removed by identifying identical download links; (2) text extraction and preprocessing, where each PDF is converted into plain text and non-textual elements such as tables and figures are removed; (3) locality-sensitive hashing (LSH) deduplication, which eliminates near-duplicate documents with overlapping content; (4) semantic chunking, where cleaned documents are segmented into passages of fewer than 256 tokens to balance retrievability and semantic integrity; and (5) embedding-based near deduplication, where dense embeddings are computed for all chunks, an ANN index is built, and pairs with cosine similarity above 0.9 are discarded. Chunks obtained are regarded as disaster management passages in DM-MTP.
In the second and third stages, given a passage and the definition of a search intent, an LLM generates information-need statement. The information statement together with the passage is used to construct a user query and its directly relevant positive passage. Prompts for query generation and relevant passage generation based on disaster management-related passages under different search intents are adapted from yin2025disastir.
@@ Appendix C Data Contamination Analysis
Following wang2023improving, we conduct test set data contamination analysis from two aspects: query overlap and passage corpus overlap.
@@
Following wang2023improving, we perform a string-matching analysis between queries in the DisastIR-Test and those in the training sets (MTP and MTT), and find no overlap between training and test queries.
@@
While sharing the same passage corpus (e.g., DBPedia, NQ, and TriviaQA all using Wikipedia) is a standard evaluation practice in information retrieval and is not regarded as contamination (wang2023improving), we adopt a stricter criterion by performing a string-matching analysis between passages in DisastIR-Test and those in MTP and MTT, finding no overlap between the training and test corpora.
@@ Appendix D Determination of $\alpha$ Value during Progressive Fine-tuning
We follow the ablation study results of moreira2024nv and set $\alpha=0.95$ for DMRetriever-596M, 4B, and 7.6B.
@@ Appendix G Detailed Reasons for Excluding Knowledge Distilled (KD) Baselines
For baselines (wang2022text; awasthy2025granite) that use knowledge distillation (KD) during fine-tuning, a cross-encoder is typically employed as the teacher model. This process introduces substantial external knowledge, as the teacher cross-encoder has a more powerful architecture. Unlike bi-encoders (our primary baselines and DMRetriever), cross-encoders process queries and documents jointly, allowing deep token-level interactions that produce highly accurate relevance scores. These fine-grained scores are then used as soft labels for the student model, offering a much richer and more informative training signal than the sparse binary hard labels available to other models.
Consequently, student model does not learn from ground-truth data alone, but is guided by teacher’s high-quality interpretation of that data.
For this reason, we exclude all KD-based baselines, which include E5-small, base, large-v2 (wang2022text), and granite-embedding-125m-english (awasthy2025granite), to ensure fair comparisons.
@@ Appendix H Information on Baseline Models and Their Implementation
We adopt baseline models evaluated in DisastIR and additionally include recent models from MTEB that were not evaluated in DisastIR. They are categorized as four scales based on model parameter sizes: small (<=109M), medium (137M - 335M), large (434M- 1.5B), and extra large (XL) (>= 4B).
Detailed specifications of the selected models are provided in Table 12, while their HuggingFace links and licenses are listed in Table 13.
@@ Appendix I Implementation Details of DMRetriever Training
For unsupervised contrastive pre-training, each model is trained for 1 epoch. We use 3 H200 GPUs for encoder-only backbones and 4 H200 GPUs for decoder-only backbones. To optimize GPU memory consumption, we adopt DeepSpeed ZeRO Stage 1 (rasley2020deepspeed) and gradient checkpointing. Gradient accumulation is set to 1, and the per-GPU batch size is maximized to increase the diversity of in-batch negatives (bge_embedding), leading to the global batch sizes reported in Table 7.
@@
Across all experiments, we use brain floating point (bfloat16) quantization and truncate input sequences to 512 tokens. The temperature value is set as 0.01. For each query, we set one positive passage and nine passages during pre-training and fine-tuning. For decoder-only backbones, we apply LoRA with rank $r=16$ and $\alpha=32$, following prior work (lee2024nv; wang2023improving; xu2024bmretriever), leaving hyperparameter tuning for future exploration.
@@
To implement the CBF method, we follow the methodology of wang2022text. Specifically, for each search intent, we use its best-performing retrieval model identified from the DisastIR benchmark as a reference (Table 4 for best-performing retrieval models). A synthetic pair is then discarded if its LLM-generated passage is not ranked within the top-2 results retrieved by this reference model.
@@
Appendix L Performance of DMRetriever-KD
To establish a fair comparison with baselines like the E5 series that employ knowledge distillation (KD) during model fine-tuning and boost performance of our smaller models, we apply KD to our fine-tuned DMRetriever-33M, 109M, and 335M models. This process yields DMRetriever-KD series (see Appendix M for KD settings). As shown in Table M, our DMRetriever-KD achieves the best average performance within their respective scales, securing the top rank across the majority of search intents.
[…]
Appendix M Knowledge Distillation Settings
We apply knowledge distillation (KD) to our DMRetriever-33M, 109M, and 335M models, creating their distilled counterparts, denoted as the DMRetriever-KD series. The distillation is performed using a specially constructed dataset, MTT-0.95-KD.
This dataset is an extension of MTT-0.95, the construction of which is detailed in Section 3.4 and Algorithm 3 (Lines 10–16). To create KD version, we use the best-performing IR model, as identified in DisastIR, for each search intent as a teacher to generate query-passage similarity scores. These scores are then added to the dataset as soft labels.
The training objective for knowledge distillation, $\mathcal{L}_{\text{KD}}$, is formulated as a weighted sum of two components. It combines the contrastive loss, $\mathcal{L}_{\text{CL}}$ (from Eq. 4), on the original hard labels with a Kullback-Leibler (KL) divergence term, $\mathcal{L}_{\text{KL}}$, for distilling the soft labels from the teacher. The balance between these two is controlled by a hyperparameter $\lambda$, as shown below: L_KD = L_CL + λL_KL In this work, we set the weighting factor $\lambda$ to 1 for all distillation experiments.
Model Scale QA QAdoc TW FC NLI STS Avg
[…]
e5-base-V2 (wang2022text) 109M 65.52 62.68 57.53 61.86 45.40 74.38 61.23
e5-small-V2 (wang2022text) 33M 65.79 62.64 60.12 61.80 46.99 74.54 61.98
[…]
granite-embedding-125m (awasthy2025granite) 125M 64.58 60.73 46.47 62.45 48.05 71.61 58.98
e5-large-V2 (wang2022text) 335M 59.92 63.07 55.37 61.85 50.69 74.69 60.93
[…]
Table 8: Performance of knowledge distillation enhanced retrieval models across multiple search intents.